# Hypersonic flow onto a large curved wedge and the dissipation of shock wave

### v1

@@ 5.3. Global solution
Define $\bar{\xi}:=\sup\Gamma$. We assume (with the aim of arriving at a contradiction) $\bar{\xi}\neq+\infty$. Then the $C^{-}$ characteristics is tangent to the wedge $\mathsf{W}$ on the point $\mathsf{W}|_{\bar{\xi}}$. Otherwise, we can extend the solution forward a little and it is contradict to $\bar{\xi}=\sup\Gamma$. So we have $\beta=\tau$ at the point $\mathsf{W}|_{\bar{\xi}}$. This indicates $\omega=\tau-\beta=0$ and the flow is vacuum. But, by (2.2.8), we have ¯$\bar{\partial}^{-}\tau\leq 0$. So $\tau|_{P^{s}}\geq\tau|_{C^{-}}\geq\tau|_{\underline{P}_{1}}\geq 0$. In view of Theorem 5.4, we can derive that $(u,v)|_{C^{-}}$ locates in the upper $(u,v)$ plane and is bounded by the curves: (1) the line $v=0$; (2) the circle $q=q|_{P^{s}}$; (3) the shock polar $G(\vec{U},\epsilon)=0$. So the states $(u,v)$ on the point $\mathsf{W}|_{\bar{\xi}}$ must locate inside the $(u,v)$ domain bounded by the shock polar $G(\vec{U},\epsilon)=0$. It indicates that the flow can not touch the vacuum along $C^{-}$. Thus we get contradiction and $\bar{\xi}=+\infty$.
In a similar way, we can show that all the $C^{+}$ characteristics issuing from the wedge $\mathsf{W}$ arrive at the shock $\mathsf{S}$. To see this, we assume to the contrary that there exists a $C^{+}$ characteristics tangent to the shock $\mathsf{S}$ on a point $\mathsf{S}|_{\bar{\xi}}$. So, the shock is degenerate on $\mathsf{S}|_{\bar{\xi}}$. On the one hand, $\mathsf{S}$ is convex for $\xi\in[\xi_{0},\bar{\xi}]$. On the other hand, (2.2.8) yields $c\bar{\partial}^{+}\alpha=-\nu\Xi\sin(2\omega)\bar{\partial}^{+}c>0$, where the hypersonic incoming flow assumption guarantee $\Xi>0$. The opposite convexity between $\mathsf{S}$ and $C^{+}$ characteristic at the tangent point $\mathsf{S}|_{\bar{\xi}}$ leads to a contradiction. ∎
Above lemma leads to ¯±$\bar{\partial}^{\pm}c<0$ for $\xi\geq\xi_{0}$. Finally, we give the lower boundedness of ¯±$\bar{\partial}^{\pm}c$. As illustrating in Figure 10, we let $P\in\Omega$ and consider ¯$\bar{\partial}^{+}c(P)$. By ¯±$\bar{\partial}^{\pm}c<0$, the terms in brackets of (2.2.9) are negative, we have
@@ 5.4. Asymptotic behaviour
By Theorem 5.7, the shock is convex and $\tau\geq 0$, so $\lim\limits_{x\rightarrow+\infty}\phi^{\prime}(x)=\inf\phi^{\prime}(x)$ exists. In the proof there, we also deduce that $\Omega$ is covered by the $C^{-}$ characteristics from the shock. Therefore, on the $(u,v)$ plane, the flow fields is covered by the contour line $F(u,v)=\mbox{constant}$ from the shock polar (2.1.2), where $F$ is the Riemann invariant for the -characteristic. Then, in view of Theorem 5.4, it suffices to show

### v2

@@ 5.3. Global solution
Define $\bar{\xi}:=\sup\Gamma$. We assume (with the aim of arriving at a contradiction) $\bar{\xi}\neq+\infty$. Then the $C^{-}$ characteristic is tangent to the wedge $\mathsf{W}$ on the point $\mathsf{W}|_{\bar{\xi}}$. Otherwise, we can extend the solution forward a little and it contradicts $\bar{\xi}=\sup\Gamma$. So we have $\beta=\tau$ at the point $\mathsf{W}|_{\bar{\xi}}$. This indicates $\omega=\tau-\beta=0$ and the flow is vacuum. But, by (2.2.8), we have ¯$\bar{\partial}^{-}\tau\leq 0$. So $\tau|_{P^{s}}\geq\tau|_{C^{-}}\geq\tau|_{\underline{P}_{1}}\geq 0$. In view of Theorem 5.4, we can derive that $(u,v)|_{C^{-}}$ locates in the upper $(u,v)$ plane and is bounded by the curves: (1) the line $v=0$; (2) the circle $q=q|_{P^{s}}$; (3) the shock polar $G(\vec{U},\epsilon)=0$. So the states $(u,v)$ on the point $\mathsf{W}|_{\bar{\xi}}$ must locate inside the $(u,v)$ domain bounded by the shock polar $G(\vec{U},\epsilon)=0$. It indicates that the flow can not touch the vacuum along $C^{-}$. Thus we get contradiction and $\bar{\xi}=+\infty$.
In a similar way, we can show that all the $C^{+}$ characteristics issuing from the wedge $\mathsf{W}$ arrive at the shock $\mathsf{S}$. To see this, we assume to the contrary that there exists a $C^{+}$ characteristic tangent to the shock $\mathsf{S}$ on a point $\mathsf{S}|_{\bar{\xi}}$. So, the shock is degenerate on $\mathsf{S}|_{\bar{\xi}}$. On the one hand, $\mathsf{S}$ is convex for $\xi\in[\xi_{0},\bar{\xi}]$. On the other hand, (2.2.8) yields $c\bar{\partial}^{+}\alpha=-\nu\Xi\sin(2\omega)\bar{\partial}^{+}c>0$, where the hypersonic incoming flow assumption guarantee $\Xi>0$. The opposing convexity of $\mathsf{S}$ and the $C^{+}$ characteristic at the tangent point $\mathsf{S}|_{\bar{\xi}}$ leads to a contradiction. ∎
Above lemma leads to ¯±$\bar{\partial}^{\pm}c<0$ for $\xi\geq\xi_{0}$. Finally, we give the lower boundedness of ¯±$\bar{\partial}^{\pm}c$. As illustrated in Figure 10, we let $P\in\Omega$ and consider ¯$\bar{\partial}^{+}c(P)$. By ¯±$\bar{\partial}^{\pm}c<0$, the terms in brackets of (2.2.9) are negative, we have
@@ 5.4. Asymptotic behaviour
By Theorem 5.7, the shock is convex and $\tau\geq 0$, so $\lim\limits_{x\rightarrow+\infty}\phi^{\prime}(x)=\inf\phi^{\prime}(x)$ exists. In the proof there, we also deduce that $\Omega$ is covered by the $C^{-}$ characteristics from the shock. Therefore, on the $(u,v)$ plane, the flow field is covered by the contour line $F(u,v)=\mbox{constant}$ from the shock polar (2.1.2), where $F$ is the Riemann invariant for the -characteristic. Then, in view of Theorem 5.4, it suffices to show